OpenStudy (anonymous):

What are the approximate solutions of 2x^2 + 9x = 8 to the nearest hundredth?

OpenStudy (cwrw238):

2x^2 + 9x - 8 = 0 you can use the quadratic formula to solve this x = [ -b +- sqrt(b^2 - 4ac)] / 2a

OpenStudy (cwrw238):

here a, b and c are the values in the general equation ax^2 + bx + c = 0

OpenStudy (cwrw238):

need any ore help?

OpenStudy (cwrw238):

*more

OpenStudy (anonymous):

yea

OpenStudy (cwrw238):

ok - 2x^2 + 9x - 8 = 0 - compare this with ax^2 + bx + c = 0 so a = 2, b = 9 and c = -8 plug these into the formula: x =[ -9 + sqrt(9^2 - 4 * 2 * -8)] / 2*2 = [ -9 + sqrt(81+64)] / 4 = -9/2 + sqrt 145 also other root is -9/2 - sqrt 145 these work out to 7.54 or -16.54 to nearest one hundredth

OpenStudy (anonymous):

then what?

OpenStudy (cwrw238):

no more those are the solutions: (7.54,-16.54)

OpenStudy (anonymous):

the answers are suppose to be one of these. x ≈ 1.22 and x ≈ 3.28 x ≈ -0.76 and x ≈ 5.26 x ≈ -1.22 and x ≈ -3.28 x ≈ -5.26 and x ≈ 0.76

OpenStudy (cwrw238):

oh - ok - i know my method is ok - i might have made a mistake in the calculations i'll re-check

OpenStudy (anonymous):

brb

OpenStudy (cwrw238):

yeah - i see my mistake its on last line it should be -9/4 + sqrt145/4 = -2.25 + 3.01 = 0.76 and -2.25 - 3.01 = -5.26 answer 0.76 and -5.26